2. After a mixture of hydrogen and nitrogen gases in a reaction vessel is allowed to attain equilibrium at 472℃, it is found to contain 7.38 atm H2, 2.46 atm N2, and 0.166 atm NH3. From these data,
a. Write the chemical equation of the reaction.
b. Calculate the equilibrium constant Kp for this reaction.
1
Expert's answer
2020-10-26T14:53:52-0400

Chemical Equation:

"N_2+3H_2\to2NH_3"


"K_P=\frac{P_{NH_3}^2}{P_{H_2}^3\times P_{N_2}}=\frac{0.166\times0.166}{2.46\times(7.38)^3}=2.78\times10^{-5}"